OpenStudy (anonymous):

Find the vertex focus and directrix of the parabola y^2+3y=-x-2

OpenStudy (phi):

I would put the equation into general form: x= -y^2 -3y - 2 do you know how to complete the square, so that you can change this into vertex form?

OpenStudy (anonymous):

No i am not sure this is where i am stuck

OpenStudy (anonymous):

can you show me what to do next? i watched the video and still confused

OpenStudy (campbell_st):

well the equation is a parabola in the form \[(y-k)^2 = 4a(x-h) \] where ((h, k) is the vertex and a is the focal length so complete the square in y by adding a value to both sides of the equation it basically looks like 

OpenStudy (phi):

this video is shorter and closer to your problem http://www.khanacademy.org/math/algebra/quadratics/completing_the_square/v/ex3-completing-the-square

OpenStudy (phi):

though in your problem, we switch the x and y 's we want to change x= -y^2 -3y - 2 to the form x = a(y-h)^2 + k (h,k) will be the vertex

OpenStudy (anonymous):

-y(y+3)-2=x

OpenStudy (anonymous):

?

OpenStudy (ranga):

y^2 + 3y = -x - 2 complete the square. Take coefficient of y term, divide by 2: 3/2 3/2 will go inside the square and you need to add (3/2)^2 to the right side: (y + 3/2)^2 = -x - 2 + (3/2)^2 = -x - 2 + 9/4 = -x - 8/4 + 9/4 = -x + 1/4 (y + 3/2)^2 - 1/4 = -x multiply both sides by -1 -(y + 3/2)^2 + 1/4 = x x = -(y + 3/2)^2 + 1/4 compare it to x = a(y - k)^2 + h a = -1; h = 1/4; k = -3/2 Vertex at (1/4, -3/2) If p is the distance from the vertex to the focus then, p = |1/4a| = 1/4 Since a is negative, this is a parabola that opens in the negative x direction. The focus will be to the left of the vertex. Subtract 1/4 from the x value of the vertex. Focus will be at (1/4 - 1/4, -3/2) = (0, -3/2) The directrix will be at distance p to the right of the vertex. (remember the parabola lways opens away from the directrix. It will never intersect with the directrix). Add 1/4 to the x value of the vertex: 1/4 + 1/4 = 1/2 The equation of the directrix is x = 1/2 (a vertical line at x = 1/2) Answers: Vertex at (1/4, -3/2) Focus at (0, -3/2) Equation of directrix: x = 1/2

OpenStudy (anonymous):

so would this parabola open to the left?

OpenStudy (ranga):

yes.